# JEL: C81, O32

### v1

@@ 2. Productivity, Capital, and Medical Research
The revival of interest in tools, such as industrial policy, to better allocate public resources to research and development is a reaction, in part, to two stylized facts (Council of Economic Advisers,, 2021, 2024; Lindsey and Hammond,, 2020; Hausmann,, 2023). First, the growth rate of research output has been constant (i.e., the left-hand side of (2.1) is constant). Second, the quantity of resources devoted to research (i.e., the factor $S_{t}$) has increased substantially. The linear specification (2.1) then implies that research productivity $\alpha_{t}$ has declined. That is, ideas are getting harder to find. For a review of the relevant literature, see Goldin et al., (2024). Detailed characterization of these facts is due to Bloom et al., (2020).
2.1. Heterogeneity
We propose an alternative reconciliation of these facts. Our hypothesis hinges on heterogeneity. If measured increases in the quantity of research are accompanied by a compositional shift toward unproductive or low-capital research, then constant growth does not necessarily imply declining productivity for capital-intensive research. That is, large increases in the quantity of low-capital-intensity or low-productivity research need not have quantitatively important impacts on growth.
To fix ideas, let the quantity $\mathcal{I}_{t}$ index the set of “research units” available at time $t$. There are two types of research units, high, $\mathsf{H}$, and low, $\mathsf{L}$. Each research unit of type $\mathsf{H}$ has a stock of research capital $S^{\mathsf{H}}_{t}$ and a productivity coefficient $\alpha^{\mathsf{H}}_{t}$. Similarly, research units of type $\mathsf{L}$ have research capital and productivity $S^{\mathsf{L}}_{t}$ and $\alpha^{\mathsf{L}}_{t}$, respectively. Consider the generalized production function
$\frac{\text{d}A_{t}}{\text{d}t}\frac{1}{A_{t}}=|\mathcal{I}^{\mathsf{H}}_{t}|\text{ }\alpha^{\mathsf{H}}_{t}S^{\mathsf{H}}_{t}+|\mathcal{I}^{\mathsf{L}}_{t}|\text{ }\alpha^{\mathsf{L}}_{t}S^{\mathsf{L}}_{t}\leavevmode\nobreak\ ,$ (2.2)
where $\mathcal{I}^{\mathsf{H}}_{t}$ and $\mathcal{I}^{\mathsf{L}}_{t}$ index the set of $\mathsf{H}$-type and $\mathsf{L}$-type research units and ||$\cdot$|| ⋅ | denotes the cardinality operator. That is, relative to the canonical specification (2.1), we entertain heterogeneity across research units in both productivity and capital-intensity.
Suppose that the quantity, productivity, and capital of $\mathsf{H}$-type units is constant over time. Thus, in each period $t$, the first term in (2.2) is given by $|\mathcal{I}^{\mathsf{H}}|\text{ }\alpha^{\mathsf{H}}S^{\mathsf{H}}$, where we have removed the time subscript. Suppose that at time $t=0$, there are no $\mathsf{L}$-type units, and that at time $t=1$, the quantity of $\mathsf{L}$-type units $\mathcal{I}^{\mathsf{L}}_{1}$ becomes non-zero. The impact of this change only affects growth through the factors $S_{1}^{\mathsf{L}}$ and $\alpha_{1}^{\mathsf{L}}$. If either quantity is negligible, relative to the contribution of the $\mathsf{H}$-type units, then there is a negligible impact on growth.
This story is closely related to a strand of the economic growth literature that argues that measured increases in the quantity of research reflect increases in the number of product varieties (Dinopoulos and Thompson,, 1998; Peretto,, 1998; Young,, 1998; Howitt,, 1999). By contrast, we examine the idea that shifts in the relative shares of established varieties can rationalize the same trends.
2.2. Medical Research
The objective of this paper is to test the compositional hypothesis described in Section 2.1. We do so in the context of medical research. We argue that in this setting we are able to decompose research units into coarse types. Moreover, for each type, institutional context allows us to infer absolute and relative productivity and capital-intensity. Measures of quantity can be recovered from existing data. An additional advantage of this setting is that medical research has an intuitive notion of output: improvement in mortality. To keep our focus on the composition of inputs—the right-hand side of (2.2)—we note that it is a well-established empirical regularity that life expectancy has risen at a steady rate since 1840 (see e.g., Oeppen and Vaupel,, 2002; Bloom et al.,, 2020). We proceed assuming that the left-hand side of (2.2) is constant, as is standard in this literature. [^9] We make the assumption that life expectancy is improving at a steady rate, given the stability of these trends for more than 150 years. It is of course unlikely that medical innovations developed in the early 2010s are clearly discernible in recent mortality data, especially given the impact of the COVID-19 pandemic.
Medical research takes many forms. We focus on two types: clinical trials, and other forms of clinical research that aim to inform the provision of medical care. Clinical trials are expensive, time-consuming, highly-regulated experiments, which have as their goal the production of causal and mortality-relevant evidence (Alsan et al.,, 2024; Wouters et al.,, 2020). The contours of the second set, other forms of clinical research, are less precise. This set includes observational studies, meta-analyses and systematic reviews, cost-effectiveness studies, etc. (Food and Drug Administration,, 2024). In contrast to clinical trials, these other study types are less capital-intensive and less productive, where productivity is defined as their efficiency in turning research inputs into changes in mortality (Jones and Podolsky,, 2015; Chavez-MacGregor and Giordano,, 2016; Kones et al.,, 2014). [^10] Of course, one can identify specific studies that deviate from our categorization. See Kones et al., (2014) for discussion of a setting in which, the authors argue, observational data are as informative as randomized controlled trial evidence. However, clinical trials remain the “gold standard” for medical research (Jones and Podolsky,, 2015). Non-randomized evidence is consistently excluded from prescribing guidelines, on the grounds that it provides evidence of lesser quality. Chavez-MacGregor and Giordano, (2016) and Kones et al., (2014) discuss hierarchies of medical evidence in more detail.
3. Identifying Clinical Trials in Publication Data
We construct a census of clinical trials. No existing dataset, used off-the-shelf, is sufficient. Although dollar-denominated expenditures are an obvious metric, available records of investment in clinical trials are self-reported, in some form, by pharmaceutical firms and are essentially impossible to validate independently. [^11] There are two frequently used sources of data on dollar-denominated expenditures. DiMasi et al., (2016) provide one set of estimates of drug research and development costs, based on confidential data obtained from 10 firms. Myers and Pauly, (2019) and Scannell et al., (2012) use research and development expenditure figures published by the Pharmaceutical Research and Manufacturers of America, a trade organization representing the industry. Administrative databases—such as ClinicalTrials.gov—aim to provide an index of trials, but suffer from widespread non-compliance with reporting requirements, with substantial variation in enforcement over time (DeVito et al.,, 2020). Proprietary databases pose two challenges: first, it is, again, difficult to validate their contents and, second, leading databases take as inputs data in public registries, thus adopting the same shifts in reporting over time (see e.g., the list of registry inputs for one frequently-used database, Cortellis,, 2024). For each of these options, there is no symmetric way to identify the quantity of other clinical research.
3.1. Publication Data
@@
Publication data provide a natural alternative. The database PubMed / MEDLINE indexes roughly $34$ million publications, a near universe of published biomedical research. [^12] See Section A.1 for further details on on our treatment of the PubMed data. Scientific publications provide an intuitive proxy for the quantity of research inputs. [^13] The existence of publication bias means that scientific publications do not perfectly capture research inputs. See Andrews and Kasy, (2019) for a discussion. In a setting such as ours—where objects of interest are trends in research over time—we need only assume that changes in publication bias year-on-year do not perfectly offset changes in research investment. Of these $34$ million records, which are clinical trials?
@@
Approaches used by researchers in social science, instead, largely draw on metadata contained in PubMed. [^14] Oostrom, (2024) and Kao et al., (2023) are two exceptions. Oostrom, (2024) uses data from a meta-analysis of published clinical trials. Kao et al., (2023) collect trial records from the proceedings of scientific conferences. Lichtenberg, (2018) and Bloom et al., (2020) collect all records assigned a “publication type”—in the PubMed indexing process—of “clinical trial.” Feldman et al., (2019) collect all records with a publication type similar to clinical trial (e.g., also including “randomized controlled trial”).
@@
On inspection, these tags are imprecise. Both approaches retrieve large numbers of records that are not, in fact, clinical trials, while excluding potentially relevant records. Figure 1 contrasts counts of published trials constructed in various ways. The light green line (‘Clinical Trial’ NLM Tag) plots the number of publications, in each calendar year, indexed with the publication type “clinical trial.” The teal line (Any NLM Tag) plots counts of records indexed with any of 18 types that are likely to include clinical trials or related medical research. [^15] On inspection, it appears as though the NLM shifted from using the tag “Clinical Trial” to “Randomized Controlled Trial”, as well as more specific tags (e.g., “Clinical Trial, Phase 1”) . Alternative approaches to identifying trials are similarly imprecise. Medical journals increasingly require trials to be registered in open registries, e.g., ClinicalTrials.gov, as a condition of publication. The blue line (Any Registry Tag) counts records that report, in their abstract, an identifier associated with one of the four largest international trial registries. Language in a record’s abstract can indicate that it might be a clinical trial if, for example, it references a “treatment group” and “control group.” The dark blue line (Any Keyword) plots the count of records with such keywords over time. We define the “potential universe” of clinical trials as those publications with any of the following: a publication type variable similar to clinical trial, a registry identifier reported in its abstract, or a keyword reported in its abstract. The purple line (Potential Universe) plots this trend over time. Figure 1 highlights that alternative approaches to data construction, in this setting, yield meaningfully different conclusions about levels, trends, and composition. See Section A.2 for further details on the construction of these series.
Figure 1. Universe of Potential Clinical Trials
@@
Notes: Figure 1 displays counts of the number of clinical trials indexed in PubMed / MEDLINE over time, constructed using alternative search strategies. The National Library of Medicine (NLM) categorizes each publication into a “pubtype.” The light green line (‘Clinical Trial’ NLM Tag) displays the number of publications in the “Clinical Trial” pubtype. The teal line (Any NLM Tag) displays the number of publication whose pubtype is an element of a set of 18 categories likely to include clinical trials. The blue line (Any Registry Tag) gives the number of publications that report, in their abstract, an identifier associated with one of the four largest international trial registries. The dark blue line (Any Keyword) indicates the number of publications whose abstract contains a keyword indicative of a clinical trial. The purple line (Potential Universe) displays the number of clinical trials in the union of the sets of publications identified with the other lines. See Section A.2 for further details.

### v2

@@ 2. Measuring Research
Although each element of this framework is simple, constructing their empirical analogues is challenging. Decades of scholarship—summarized in Griliches, (1990, 1994) and, more recently, Bryan and Williams, (2021)—have grappled with alternative ways of measuring outputs, $A_{t}$. [^17] Pakes and Griliches, (1980) consider this challenge and employ slightly different notation than we use here. They focus on the quantification of an output variable $K$, which is the “level of economically valuable technological knowledge.” In this paper, our focus is on the comparable challenges associated with quantifying the resources that go into their production, $S_{t}$.
An ideal measure of resources $S_{t}$ should capture the total quantity of economic resources devoted to increasing the stock of useful knowledge. Not all economic activities related to scientific research are equally relevant. Many forms of scientific output have as their goal the synthesis or diffusion of existing information. In medicine, for example, literature reviews and meta-analyses allow physicians to rapidly review the empirical evidence supporting particular treatment options. [^18] For one discussion of the appropriate interpretation of meta-analyses in clinical practice, see Thompson and Pocock, (1991). Medical guidelines, produced by entities like the American Heart Association (e.g., Smith Jr. et al.,, 2001), consolidate these findings into consensus documents that may be viewed as defining the “standard of care” (Moffett and Moore,, 2011). These types of literature syntheses are, anecdotally, of increasing importance, as health technologies previously available only in a small number of countries are made available to new patient populations (Ioannidis et al.,, 2013; Cockburn et al.,, 2016).
It is important to be clear about the distinction we are making between these forms of research. This paper, and productivity calculations more generally, are interested in those forms of research that generate new information. These are the elements of the quantity $S_{t}$. From a welfare perspective, there are many other types of investment that are important, including these types of literature reviews that aid in the diffusion of information. These other investments may enable the “resources related by productivity growth . . . [to] find adequate employment,” thus increasing aggregate welfare, but are not themselves contributing to productivity growth (Griliches,, 1994).
Even if, conceptually, it is clear that certain economic activities are more relevant to the quantity $S_{t}$ than others, available indicators of research—such as scientific publications and patents—may make it challenging to accommodate this heterogeneity empirically. This task is somewhat simpler in industries often studied by economists (see, e.g., the discussion of manufacturing firms in Syverson,, 2011), where dollar-denominated expenditures and price indices are available. In many settings central to economic growth, however, research occurs in university-based laboratories that do not maintain detailed financial ledgers, [^19] The UMETRICS data are one exception. See Babina et al., (2023) for an example of a study employing financial data from a university research setting. in countries where high-quality data may not be collected or made available, or in industries where confidentiality restrictions limit the public release of dollar-denominated expenditure records. Consequently, measures of research must rely on counts of researchers (who differ in skill), publications (which differ in relevance and resource-intensity), and administrative databases (which differ in completeness).
The objective of this paper is to address these conceptual and empirical challenges in the context of medical research. Measuring the outputs of the production function (2.1) is, in some sense, more straightforward in medicine than in other sectors. Medical innovation appears coarsely in records maintained by regulators (e.g., counts of “new molecular entities” approved by the U.S. Food and Drug Administration) and, more directly, in measures of life expectancy and life-years saved from disease. [^20] Cutler and McClellan, (2001) observe that technological change has accounted for the bulk of increases in health spending over time, but that health has improved alongside these increases. In a series of case studies, they argue that the benefits of technological progress exceed their costs. See, also, Almond et al., (2010), who estimate that the returns to medical spending are positive on the margin. As this paper focuses on quantification of inputs $S_{t}$, we note only that it is a well-established empirical regularity that improvements in life expectancy, on many margins, appear to be “stagnating” (Mehta et al.,, 2020; Ho and Hendi,, 2018). In drawing conclusions about research productivity, we will follow Bloom et al., (2020) and assume that the left-hand side of (2.1) is constant. [^21] It is of course unlikely that medical innovations developed in the early 2010s are clearly discernible in recent mortality data, especially given the impact of the COVID-19 pandemic.
Our task, then, is to identify a measure for inputs $S_{t}$ that best explains potential shifts in this output measure. Institutional features of medicine are helpful in doing so: medical decision-makers enforce a clear hierarchy of statistical evidence (Jones and Podolsky,, 2015). As guidelines to medical journal authors and editors from the American Medical Association state, “[p]rospective studies generally yield more reliable conclusions than retrospective studies. Cause and effect cannot be established by observational studies; consequently, study findings from observational studies should not be presented using causal language.” [^22] In 2024, the Journal of the American Medical Association—the “trunk journal” (see Section 4) of the American Medical Association and among the highest ranked journals in medicine—adopted a framework to allow studies other than randomized experiments to use this type of “causal” language. In doing so, editors of JAMA noted that “well-conducted randomized clinical trials remain the preferred approach” for the production of this type of evidence (Dahabreh and Bibbins-Domingo,, 2024). In decisions about the approval of new medicines (Food and Drug Administration,, 2024), the terms of insurance coverage (Choudhry,, 2017), and the appropriateness of a particular prescription (Martí-Carvajal,, 2020; Alsan et al.,, 2024), decision-makers defer to evidence generated via clinical trial. Other forms of medical evidence, generated through observational studies and surveys of the literature, receive less weight. Although detailed cost information on medical research is not available, available evidence suggests that these other forms of medical research are lower cost [^23] Recent estimates suggest that the average clinical trial may cost upwards of \$20 million dollars (Martin et al.,, 2017). Available estimates suggest that the cost of completing a systematic review or meta-analysis can reach \$150,000, where the bulk of the cost stems from the quantity of labor required to identify trials in publication data (Michelson and Reuter,, 2019). and involve fewer regulatory hurdles (Congressional Budget Office,, 2021). [^24] For example, while the approval of an institutional review board is necessary for any randomized clinical trial to proceed, many observational studies are exempt from this type of review. See guidance to researchers at Stanford University here: https://stanfordmedicine.box.com/shared/static/ase46ixmatdeahj7qdfgp4ydmy5qc3ei.pdf.
Thus, we treat clinical trials as the appropriate measure of research inputs $S_{t}$. Clinical trials are discrete, regulated experiments. Research proceeds according to a pre-defined experimental protocol under the supervision of a regulator and across various phases. Trial phases differ in their scale (number of participants and duration) and, thus, in their cost (Food and Drug Administration,, 2024). Standardized measures (e.g., statistical objects like p-values) are recorded at various intervals. Trials are sufficiently expensive (Congressional Budget Office,, 2021) as to reflect “real” and substantial investment, and decisions to progress a study across phases (e.g., to move from a small Phase II study to a large Phase III study) capture decisions to pursue continued, directed research in an area of study (for an application of this insight, see Cunningham et al.,, 2021).
As we approach this exercise with the ultimate goal of making progress on the question of research productivity, it is useful to note an additional feature about clinical trials, which suggests that sharp increases in quantities year-on-year (as documented in Bloom et al., (2020)) are unlikely. Not only are clinical trials sufficiently expensive as to require access to a source of financing, but trials also require an often scarce “input:” eligible patients. Malani and Philipson, (2019) collect empirical evidence of this type of “subject market effect,” and Alsan et al., (2024) collect anecdotal evidence on the difficulties of patient recruitment. [^25] The popular press has described this as a conundrum for cancer research: “too many drug trials, too few patients” (Kolata,, 2017).
3. Identifying Clinical Trials
We construct a new census of clinical trials disclosed in scientific publications. We begin, in Section 3.1, by detailing the deficiencies of existing, or alternative, approaches to measuring this form of medical research. In Sections 3.2 and 3.3 we propose a method, based on fine-tuning a large language model, for identifying relevant records at scale and with high fidelity. The performance of the method is assessed in Section 3.4.
3.1. Existing Approaches
Although many datasets report information on clinical trials—and have been increasingly used by economists to measure innovation (see e.g., Budish et al.,, 2015; Cunningham et al.,, 2021)—no existing dataset, used off-the-shelf, yields accurate, transparent quantity measures for medical research in aggregate.
3.1.1. Dollar Denominated Data
Although disaggregated, dollar-denominated trial costs are a natural measure, available records of investment in clinical trials are self-reported, in some form, by pharmaceutical firms and are essentially impossible to validate independently (for a discussion, see Morten et al.,, 2024).
There are, to our knowledge, three such sources of data that have been used in research. DiMasi et al., (2003, 2016) produce one widely cited (and contested, see Light and Warburton, (2005)) set of estimates of drug research and development costs, based on confidential data obtained from 10 firms. Myers and Pauly, (2019) and Scannell et al., (2012) collect aggregate estimates of annual research and development expenditures published by the Pharmaceutical Research and Manufacturers of America, a trade organization representing the industry. Both sets of estimates are often characterized as an upper bound (Morten et al.,, 2024) and cannot be produced independently by external researchers. Wouters et al., (2020) generates estimates of research and development costs, drawing on the information available in financial filings collected by regulators in the United States. Although these estimates are important and informative in a context that lacks precise measures of costs, expenditures can be defined in this way for only roughly 18 percent of the new drugs approved by the FDA between 2009 and 2018.
These estimates, then, are not sufficient to answer the question this paper sets out to address. Even if dollar-denominated measures were available, it is not obvious how these measures should be scaled to allow for comparison across time periods. See Berndt and Cockburn, (2013) for a discussion of the difficulties associated with constructing a price index for clinical trial research. [^26] Berndt and Cockburn, (2013) construct a preliminary price index using data on trial budgets from a private firm. This firm no longer discloses data to academic researchers.
3.1.2. Trial Registries and Proprietary Databases
Nearly all clinical trials are conducted under the auspices of a regulator that requires both pre-registration and results disclosure. In the United States, the Food and Drug Administration Amendments Act (FDAAA), as passed in 2007 and revised in a ”Final Rule” that took effect in 2017, imposes these requirements on sponsors of regulated trials. Nominally, effected sponsors are required to register their studies on “ClinicalTrials.gov.” For nearly all studies, sponsors must also report the results of a trial within a designated period after completion. Moreover, increasingly, medical journals require registration as a pre-condition of publication (Laine et al.,, 2007).
In practice, there is widespread noncompliance with these regulations and requirements. Deborah Zarin, then-director of ClinicalTrials.gov, notes that “funders, sponsors, and institutional review boards continue to allow unregistered trials or trials with late registration to be conducted, and some journals continue to allow the results of such trials to potentially be published” (Zarin et al.,, 2017). Noncompliance with federal reporting requirements is similarly widespread. In a series of papers assessing compliance over time DeVito et al., (2020) and DeVito and Goldacre, (2021) find that roughly 60 percent of trials registered on ClinicalTrials.gov fail to report required data elements at the time of their reporting deadline. [^27] The FDA is empowered, by statute, to impose fines of up \$10,000 for each day that required disclosures are late. To date, no fines have been collected, but the agency now sends notices of noncompliance to delinquent sponsors (Stephenson,, 2021). Notices are available here: https://www.fda.gov/science-research/fdas-role-clinicaltrialsgov-information/clinicaltrials gov-notices-noncompliance-and-civil-money-penalty-actions
From the perspective of our study, there are three issues with the use of registry data. First, requirements for study registration have changed over time. Table B.1 lists five changes between 2000 and 2020 that altered the set of studies that should appear in a complete version of the database. Second, sponsors engage in considerable “back-filling” of past studies. Zarin et al., (2017) find that a large proportion of trials are registered after the trial start date. [^28] Often back-filling relates to changes in organizational disclosure policy. In a version of ClinicalTrials.gov current through 31 May 2023, Boehringer Ingelheim—the largest privately held pharmaceutical company in the world—appears to have posted 804 studies in 2014, only 72 of which were started in 2014. In 2013, Boehringer registered 101 studies. In the fifteen preceding years, since the creation of the database, Boehringer posted 872 records in total. We selected this case study based on a discussion in Zarin et al., (2017). Third, not all records in ClinicalTrials.gov and similar registries are, in fact, clinical trials (Tse et al.,, 2018). As a result, it is difficult to determine whether changes in the number of records reflect actual growth in scientific research or a change in the propensity to register studies (trials and non-trial studies).
In Appendix B, we illustrate the difficulty of recovering accurate measures of trends in the quantity of clinical trials from this database. We show that small changes in data construction yield meaningfully different estimates of trial levels and that many records are not in fact clinical trials. Moreover, due to widespread data gaps resulting from noncompliance in reporting of results and study attributes, there is little reliable information about characteristics of registered studies.
Other databases, such as the PDQ Cancer Information database maintained by the U.S. National Cancer Institute and used in papers including Budish et al., (2015), are disease-specific and, thus, insufficiently general for our purposes. A small number of firms (notably, GlaxoSmithKline) have disclosed all clinical trials over a longer period of time. [^29] See GlaxoSmithKline’s registry here: https://www.gsk-studyregister.com/ Of course, records from a single firm are not informative about aggregate trends, especially because many investigational drugs change hands frequently during the course of development (Cunningham et al.,, 2021).
Researchers, instead, often draw on proprietary databases. Although these databases are sufficient for certain research questions—for example, related to changes in the ownership of investigational drugs (Cunningham et al.,, 2021, using Pharmaprojects data) or on the link between insurance policy and drug development (Agha et al.,, 2022, using Cortellis data)—they are unsuitable for our particular productivity measurement exercise. Our aim is to construct a census that allows for comparison of quantities across points in time. First, we are unable to determine, with specificity, how records are identified and catalogued based on available documentation. Second, these databases take public registries as inputs and, thus, risk adopting the same shifts in reporting over time (see e.g., the list of registry inputs for one frequently-used database, Cortellis,, 2024).
3.1.3. Publication Data
@@
Publication data provide a natural alternative. [^30] In Section 4, we discuss the two natural concerns with measures constructed from publication data—publication bias and shifts in reporting norms over time. Neither is likely to affect the interpretation of trends in our data. The database PubMed / MEDLINE indexes roughly $34$ million publications, a near universe of published biomedical research. [^31] See Section A.1 for further details on on our treatment of the PubMed data. It is reasonable to expect that the corpus of scientific publications contains a census of research. The challenge is in extracting measures of research from this unstructured text. Of these $34$ million records, which are clinical trials?
@@
Approaches used by researchers in social science, instead, largely draw on metadata contained in PubMed. [^32] Oostrom, (2024) and Kao et al., (2023) are two exceptions. Oostrom, (2024) uses data from a meta-analysis of published clinical trials. Kao et al., (2023) collect trial records from the proceedings of scientific conferences. Lichtenberg, (2018) and Bloom et al., (2020) collect all records assigned a “publication type”—in the PubMed indexing process—of “clinical trial.” Feldman et al., (2019) collect all records with a publication type similar to clinical trial (e.g., also including “randomized controlled trial”). On inspection, these tags are imprecise. Both approaches retrieve large numbers of records that are not, in fact, clinical trials, while excluding potentially relevant records.
@@
Figure 3.1 contrasts counts of published trials constructed in various ways. The light green line (‘Clinical Trial’ NLM Tag) plots the number of publications, in each calendar year, indexed with the publication type “clinical trial.” This light green line is the measure reported in Bloom et al., (2020). The teal line (Any NLM Tag) plots counts of records indexed with any of 18 types that are likely to include clinical trials or related medical research. [^33] On inspection, it appears as though the NLM shifted from primarily using the tag “Clinical Trial” to “Randomized Controlled Trial” and added more specific tags (e.g., “Clinical Trial, Phase 1”) . There are 19,937 records tagged with “Clinical Trial” in 2005 and 7,518 with this same tag in 2006. Alternative approaches to identifying trials are similarly imprecise. Medical journals increasingly require trials to be registered in open registries, e.g., ClinicalTrials.gov, as a condition of publication. The blue line (Any Registry Tag) counts records that report, in their abstract, an identifier associated with one of the four largest international trial registries. Language in a record’s abstract can indicate that it might be a clinical trial if, for example, it references a “treatment group” and “control group.” The dark blue line (Any Keyword) plots the count of records with such keywords over time.
We define the “potential universe” of clinical trials as those publications with any of the following: a publication type variable similar to clinical trial, a registry identifier reported in its abstract, or a keyword reported in its abstract. The purple line (Potential Universe) plots this trend over time.
Figure 3.1 highlights that alternative approaches to data construction, in this setting, yield meaningfully different conclusions about levels, trends, and composition. That few of these series suggest “smooth” trends in the quantity of trials indicates, also, the extent to which the quality of data can affect these estimates. See Section A.2 for further details on the construction of these series.
Given these difficulties, we describe a procedure, below, to recover a sample of clinical trials from this corpus of text in a manner that is consistent and transparent.
Figure 3.1. Universe of Potential Clinical Trials
@@
Notes: Figure 3.1 displays counts of the number of clinical trials indexed in PubMed / MEDLINE over time, constructed using alternative search strategies. The National Library of Medicine (NLM) categorizes each publication into a “pubtype.” The light green line (‘Clinical Trial’ NLM Tag) displays the number of publications in the “Clinical Trial” pubtype. The teal line (Any NLM Tag) displays the number of publication whose pubtype is an element of a set of 18 categories likely to include clinical trials. The blue line (Any Registry Tag) gives the number of publications that report, in their abstract, an identifier associated with one of the four largest international trial registries. The dark blue line (Any Keyword) indicates the number of publications whose abstract contains a keyword indicative of a clinical trial. The purple line (Potential Universe) displays the number of clinical trials in the union of the sets of publications identified with the other lines. See Section A.2 for further details.